# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v2

@@ II-F Fault Characteristics Metrics
To characterize fault networks within each dataset, we measure a set of geometric and topological descriptors. These include Length, Curvature, Sinuosity, Segments, and Stepover Density, which correspond to the statistics reported in Table III.
Length ($L$) [91] is the total length of all fault traces, computed as
@@
Curvature ($\kappa$) [92] quantifies the local bending of fault traces. At each arc-length position $s$, curvature is defined as
@@
Sinuosity ($S$) [93] measures how tortuous a fault is, defined as the ratio of its length to the straight-line distance between its endpoints:
@@
Segments ($N_{\text{seg}}$) [94] is the total number of discrete fault segments observed in the network.
Stepover Density ($D_{\text{stepover}}$) [95] measures the relative occurrence of stepovers, normalized by fault length:
@@ III Fault Delineation Datasets
Among all $74$ datasets used for fault delineation [96], only $4$ field datasets (LANDMASS [97, 23, 98, 99, 100, 101, 102, 103], GSB [59, 104, 105], Thebe [105, 106], CRACKS [107]) and $4$ synthetic datasets (FaultSeg3D [15], Bi’s 3D synthetic [108], Wu’s 2D SR [109], Pochet’s 2D synthetic [110]) open-sourced both seismic data and labels. The low ratio of open-source labeled field data hinders the creation of benchmarks for training and evaluation of models.
@@
Besides the lack of public availability, the different characteristics of the datasets pose challenges to the development of generalizable DL models. Specifically, LANDMASS contains image-level fault labels that cannot be used to numerically evaluate the delineation of pixel-wise faults. GSB contains pixel-wise fault labels annotated on only $5$ crossline sections, which limits the evaluation scalability on large test data. Additionally, it is challenging to achieve generalization using only a small number of labels for finetuning [64]. In contrast, Thebe provides a large amount of pixel-level geophysicist labels across $1803$ crossline sections. CRACKS provides fault labels of varying quality collected from a group of interpreters, including a geophysicist expert, across $400$ inline sections.
@@
Among the four publicly available datasets, we select Thebe and CRACKS considering the model development and evaluation at the pixel level. All the seismic sections in Pochet’s 2D synthetic contain only one straight fault crossing the entire section, presenting less diversity in angle and density compared to the faults in FaultSeg3D. Both Bi’s 3D synthetic and Wu’s 2D SR [109] originate from FaultSeg3D using the same synthesizing workflow. Thus, we use FaultSeg3D as a reference synthetic dataset with a diverse set of faults. Below we describe and compare the three datasets used for our benchmarking study in detail.
@@
FaultSeg3D is a 3D synthetic dataset with $220$ volumes each with dimensions of $128\times 128\times 128$ [15] . In order to better approximate realistic conditions, the authors added background noise estimated from real seismic volumes. The sampling rate and the frequency of the synthetic data vary across the volume to improve the diversity of the data. An example volume is shown in Fig. 3(a).
@@
CRACKS is an open-source dataset with diverse faults delineated across $400$ inline sections of the Netherlands F3 Block [107, 111]. The authors in [24] open-sourced a fully-annotated 3D geological volume of the Netherlands F3 Block for training different models and comparing the performance with objective metrics. Thus, this volume is one of the most extensively studied geographical zones for developing DL-assisted seismic interpretation frameworks [16, 62, 64, 36, 112, 63, 34, 66, 26, 27, 113, 114, 115, 116, 67, 117, 28, 118, 119, 103, 102, 120, 121, 122, 24, 58, 59, 123, 60, 124, 125, 126, 127]. The diverse fault features in the F3 block, including major versus minor faults and varying orientations, make it an excellent seismic dataset to train and evaluate fault delineation models [128, 129]. However, the annotations in [24] do not provide pixel-wise fault labels. Thus, CRACKS open-sources fully hand-labeled fault annotations by a group of $32$ interpreters with varying degrees of expertise and a domain expert geophysicist. This dataset not only establishes a standardized benchmark for objective comparison but also can be used to investigate the impact of multiple sets of labels with varying quality. Three sets of fault annotations are used to investigate the impact of training labels with varying quality, including expert labels and two more sets of lower quality labels from two other annotators. Fig. 3(b) shows an example inline section with expert fault labels from CRACKS. CRACKS and FaultSeg3D share geological similarity in the seismic sections in addition to the similar label density.
@@
Thebe is taken from a seismic survey called Thebe Gas Field in the Exmouth Plateau of the Carnarvan Basin on the NW shelf of Australia [25, 105]. The dataset contains $1803$ labeled crossline sections of size $1737\times 3174$, making it the largest publicly-available field dataset. The seismic intensity of this dataset exhibits a low variation/standard deviation, which can be observed from the low contrast in Fig. 3(c).
@@ IV-A2 Training and Test Splits
To ensure meaningful evaluations, we adopt a consistent splitting strategy across the considered datasets. Specifically, we maximize diversity in the test set while ensuring no overlap with the training data. For each dataset, we select spatially distinct subsets to prevent redundancies and simulate deployment conditions on previously unseen segments. In CRACKS, which consists of 400 contiguous inlines, we designate the first 30 and last 30 sections as the test set, totaling 60 sections. The remaining 340 central sections are used for training. This split captures geological variability across the volume while maintaining spatial separation between training and test data. For Thebe, we use 400 sections for training and reserve 100 sections from other parts of the volume for testing, following a similar diversity-maximizing approach. For FaultSeg3D, we follow the established setup from [15], using 200 synthetic volumes for training and 20 distinct volumes for testing.
@@ IV-B Model setup
We consider eight segmentation models commonly used in the seismic literature [105, 109, 19, 130] in our fault delineation experiments, using the following nomenclature:
  • deeplab: Deeplab [131] architecture with a Resnet50 [85] backbone
  • deeplab-m:Deeplab architecture with a Mobilenet [132] backbone
  • hed: Hollistically-nested edge detection model [133]
  • rcf: Richer convolutional features [134]
  • unet: Unet [45] architecture with a Resnet50 backbone
  • unet++: Unet++ [135] architecture with a Resnet50 backbone
  • unet-b: Original Unet architecture as presented in [45]
  • segformer: Transformer-based model introduced in [53]
@@ IV-C1 Post-processing for Model Predictions
Common practices of evaluating deep fault delineation networks involve two steps [64, 22, 106, 105]: $(i)$ thresholding the network predictions to obtain binary outputs, and $(ii)$ comparing the binary outcomes with the fault test labels using a metric. The test labels are processed with the same morphological operations as the training faults in order to achieve consistent fault thickness at the input and output of a model. Consequently, the thresholding in step $(i)$ needs to be adaptive to the model and the data accordingly, followed by the same morphological operations for numerical evaluation. We compute the optimal threshold for the dataset using the Optimal Dataset Scale (ODS) metric [136]. For a model, its optimal threshold is computed using the training set, which is then applied to binarize the predictions on the test data, followed by the same morphological operation for evaluation.
@@ IV-C2 Performance Metrics
As mentioned in Section II-E, pixel-based and distance-based metrics capture different aspects of prediction quality, and each can fail to fully reflect the structural accuracy of the predicted faults, making the evaluation of fault delineation methods an inherently challenging task. Our benchmark study thus adopts a holistic approach to assess prediction quality by considering a combination of multiple metrics alongside subjective inspection. In this study we choose to report one pixel-based metric, Dice Coefficient (defined in Eq. (1)) and two distance-based metrics: BCD and the modified Hausdorff (defined in Eq. 3, and Eq. 2, respectively). These metrics have not only been extensively used in the seismic literature for model performance assessment [15, 39, 137], but allow for two different evaluation axes: pixel overlap-based and structure-based.
@@ IV-C3 Fault Characteristic Comparisons
While the individual metrics described in Section II-F quantify dataset characteristics in isolation, comparative metrics evaluate the similarity between predicted and ground-truth faults. We describe below the metrics we use in Table IV to extend the individual descriptors into differences, ratios, and statistical comparisons.
@@ V-A2 Domain Shift and Joint Training
As an additional baseline, experiments on joint training on all possible combinations of datasets in a single training round are conducted using the unet model with a ResNet50 backbone. The results for all unet experiments, including these joint training configurations, are shown in Table I. The results further showcase that features learned from CRACKS and FaultSeg3D data align together, but these learned features are not easily transferable to Thebe. On the other hand, Thebe may act as a regularizer when paired with a large dataset like FaultSeg3D, hurting performance on seen datasets (compared to training them individually) but providing modest generalization to unseen ones: performance on the unseen dataset for this setting (i.e. CRACKS) does not drop as badly as that of the unseen dataset for other joint settings (e.g. FaultSeg3D + CRACKS).
TABLE I: Performance Metrics (DICE, BCD, Hausdorff) for different training schemes on Unet. All values are rounded to three decimals. Schemes are ranked as best, second best or worst based on satisfying at least two of the three metric criteria, to account for cases where metrics disagree. (blue for highest, pink for second, and gray for the worst.)

### v3

@@ II-F Fault Characteristics Metrics
To characterize fault networks within each dataset, we measure a set of geometric and topological descriptors. These include Length, Curvature, Sinuosity, Segments, and Stepover Density, which correspond to the statistics reported in Table VI.
Length ($L$) [kim2005relationship] is the total length of all fault traces, computed as
@@
Curvature ($\kappa$) [rutter2018geometry] quantifies the local bending of fault traces. At each arc-length position $s$, curvature is defined as
@@
Sinuosity ($S$) [hanson2004style] measures how tortuous a fault is, defined as the ratio of its length to the straight-line distance between its endpoints:
@@
Segments ($N_{\text{seg}}$) [schwartz1989fault] is the total number of discrete fault segments observed in the network.
Stepover Density ($D_{\text{stepover}}$) [manighetti2021fault] measures the relative occurrence of stepovers, normalized by fault length:
@@
TABLE I: Comparison of Netherlands F3 and Thebe Seismic Datasets
Parameter Netherlands F3 (North Sea) [terranubis_f3_demo_2023] Thebe (Exmouth Plateau, Australia) [bailey2013geological]
Location Offshore Netherlands, Block F3 NW Shelf of Australia, Carnarvon Basin
Year of Acquisition 1987 2007
Area Coverage $\sim$387 km2 $\sim$1200 km2
Inline × Crossline 651 × 951 $\sim$3174 × $\sim$1803
Bin Size 25 m × 25 m $\sim$25 m × $\sim$25 m (assumed)
Record Length / Sampling 0–1.848 s TWT, 4 ms $\sim$0–4.5 s TWT, 2–4 ms (typical)
Original Purpose Jurassic/Cretaceous hydrocarbon exploration Triassic gas exploration
Public Wells / Logs 4 wells with logs (sonic/gamma; 2 with density) 2 wells (Thebe-1, Thebe-2) with gas discovery
Data Format 3D post-stack time migrated (SEG-Y, OpenDtect) 3D post-stack time migrated (SEG-Y, fault-labeled)
Structural Setting Shallow deltaic shelf, minor faults, salt dome Rifted margin, rotated fault blocks
Main Stratigraphy Miocene–Pliocene clinoforms; deeper Jurassic/Cretaceous Triassic Mungaroo Formation; Jurassic–Cretaceous seal
Reservoir Target Jurassic sandstones; shallow biogenic gas pockets Triassic fluvial sands (Mungaroo Formation)
Data Availability Fully open (TNO/dGB/OpendTect) Public fault-annotated subset; full survey open-file
Key Features Clinoforms, shallow gas, polygonal faults, salt dome Fault block trap, flat spots, complex fault network
Hydrocarbons Shallow biogenic gas (non-commercial) Confirmed dry gas field ($\sim$2–3 Tcf)
Source / Seal Biogenic gas; intraformational shale; Zechstein salt Mungaroo source/reservoir; Muderong Shale seal
@@ III Fault Delineation Datasets
Figure 4: Charaterization of the three considered datasets using fault-oriented metrics.
@@
Among all $74$ datasets used for fault delineation [an2023current], only $4$ field datasets (LANDMASS [alaudah2015curvelet, alaudah2018structure, alfarraj2018multiresolution, long2018comparative, long2015characterization, alaudah2018learning, aribido2020self, aribido2021self], GSB [di2019improving, di2018seismic, an2021deep], Thebe [an2021deep, an2020overlap], CRACKS [prabhushankar2024cracks]) and $4$ synthetic datasets (FaultSeg3D [wu2019faultSeg], Bi’s 3D synthetic [bi2021deep], Wu’s 2D SR [lin2022automatic], Pochet’s 2D synthetic [pochet2018seismic]) open-sourced both seismic data and labels. The low ratio of open-source labeled field data hinders the creation of benchmarks for training and evaluation of models.
@@
Besides the lack of public availability, the different characteristics of the datasets pose challenges to the development of generalizable DL models. Specifically, LANDMASS contains image-level fault labels that cannot be used to numerically evaluate the delineation of pixel-wise faults. GSB contains pixel-wise fault labels annotated on only $5$ crossline sections, which limits the evaluation scalability on large test data. Additionally, it is challenging to achieve generalization using only a small number of labels for finetuning [mustafa2024visual]. In contrast, Thebe provides a large amount of pixel-level geophysicist labels across $1803$ crossline sections. CRACKS provides fault labels of varying quality collected from a group of interpreters, including a geophysicist expert, across $400$ inline sections.
@@
Among the four publicly available datasets, we select Thebe and CRACKS considering the model development and evaluation at the pixel level. All the seismic sections in Pochet’s 2D synthetic contain only one straight fault crossing the entire section, presenting less diversity in angle and density compared to the faults in FaultSeg3D. Both Bi’s 3D synthetic and Wu’s 2D SR [lin2022automatic] originate from FaultSeg3D using the same synthesizing workflow. Thus, we use FaultSeg3D as a reference synthetic dataset with a diverse set of faults. We showcase the acquisition and geological properties of CRACKS and Thebe in Table I.
Given that proprietary seismic data is, by definition, not publicly accessible, our study intentionally focuses on open-source datasets to ensure reproducibility and transparency. This choice aligns with established benchmarking practices in machine learning, where the goal is to provide reproducible and extensible frameworks rather than case-specific proprietary analyses. Despite being open-source, the selected datasets span markedly different geological, statistical, and labeling characteristics, including synthetic versus real data, varying signal-to-noise ratios, and both expert and crowdsourced interpretations. These variations provide meaningful heterogeneity for assessing domain shift and model generalization. To illustrate this diversity, Figure 4 presents a comparative visualization of the three datasets using the geometric and topological fault metrics introduced in Section II-F. These metrics quantitatively capture differences in fault density, continuity, and complexity, reinforcing that the chosen datasets represent heterogeneous fault conditions suitable for benchmarking generalization. Below we describe and compare the three datasets used for our benchmarking study in detail.
@@
FaultSeg3D is a 3D synthetic dataset with $220$ volumes each with dimensions of $128\times 128\times 128$ [wu2019faultSeg] . In order to better approximate realistic conditions, the authors added background noise estimated from real seismic volumes. The sampling rate and the frequency of the synthetic data vary across the volume to improve the diversity of the data. An example volume is shown in Fig. 3(a).
@@
CRACKS is an open-source dataset with diverse faults delineated across $400$ inline sections of the Netherlands F3 Block [prabhushankar2024cracks, dgb1987netherlands]. The authors in [alaudah2019machine] open-sourced a fully-annotated 3D geological volume of the Netherlands F3 Block for training different models and comparing the performance with objective metrics. Thus, this volume is one of the most extensively studied geographical zones for developing DL-assisted seismic interpretation frameworks [chowdhury2025unified, benkert2024effective, mustafa2024visual, quesada2024benchmarking, mustafa2024explainable, mustafa2023active, kokilepersaud2022volumetric, benkert2022example, benkert2022reliable, benkert2023samples, zhou2023perceptual, chowdhury2023counterfactual, shafiq2022novel, iqbal2023blind, mustafa2021joint, mustafa2021comparative, mustafa2021man, benkert2021explaining, benkert2021explainable, aribido2021self, aribido2020self, mustafa2020spatiotemporal, mustafa2020joint, soliman2020s, alaudah2019machine, di2019developing, di2019improving, alfarraj2019semisupervised, di2019semi, di2019reflector, alaudah2019facies, mustafa2019estimation, alfarraj2019semi]. The diverse fault features in the F3 block, including major versus minor faults and varying orientations, make it an excellent seismic dataset to train and evaluate fault delineation models [ishak2018application, safari2023structural]. However, the annotations in [alaudah2019machine] do not provide pixel-wise fault labels. Thus, CRACKS open-sources fully hand-labeled fault annotations by a group of $32$ interpreters with varying degrees of expertise and a domain expert geophysicist. This dataset not only establishes a standardized benchmark for objective comparison but also can be used to investigate the impact of multiple sets of labels with varying quality. Three sets of fault annotations are used to investigate the impact of training labels with varying quality, including expert labels and two more sets of lower quality labels from two other annotators. Fig. 3(b) shows an example inline section with expert fault labels from CRACKS. CRACKS and FaultSeg3D share geological similarity in the seismic sections in addition to the similar label density.
@@
Thebe is taken from a seismic survey called Thebe Gas Field in the Exmouth Plateau of the Carnarvan Basin on the NW shelf of Australia [AN2021107219, an2021deep]. The dataset contains $1803$ labeled crossline sections of size $1737\times 3174$, making it the largest publicly-available field dataset. The seismic intensity of this dataset exhibits a low variation/standard deviation, which can be observed from the low contrast in Fig. 3(c).
@@ IV-A2 Training and Test Splits
To ensure meaningful evaluations, we adopt a consistent splitting strategy across the considered datasets. Specifically, we maximize diversity in the test set while ensuring no overlap with the training data. For each dataset, we select spatially distinct subsets to prevent redundancies and simulate deployment conditions on previously unseen segments. In CRACKS, which consists of 400 contiguous inlines, we designate the first 30 and last 30 sections as the test set, totaling 60 sections. The remaining 340 central sections are used for training. This split captures geological variability across the volume while maintaining spatial separation between training and test data. For Thebe, we use 400 sections for training and reserve 100 sections from other parts of the volume for testing, following a similar diversity-maximizing approach. For FaultSeg3D, we follow the established setup from [wu2019faultSeg], using 200 synthetic volumes for training and 20 distinct volumes for testing.
@@ IV-B Model setup
We consider eight segmentation architectures that are among the most widely used in the seismic interpretation literature [an2021deep, lin2022automatic, rs16050922, 10042020]. This selection was designed to cover the set of architectures commonly adopted for fault delineation, spanning both classical convolutional models and more recent transformer-based variants. Specifically, we include:
  • deeplab: Deeplab [chen2018encoder] architecture with a Resnet50 [he2016deep] backbone
  • deeplab-m:Deeplab architecture with a Mobilenet [sandler2018mobilenetv2] backbone
  • hed: Hollistically-nested edge detection model [xie2015holistically]
  • rcf: Richer convolutional features [liu2017richer]
  • unet: Unet [ronneberger2015u] architecture with a Resnet50 backbone
  • unet++: Unet++ [zhou2018unet++] architecture with a Resnet50 backbone
  • unet-b: Original Unet architecture as presented in [ronneberger2015u]
  • segformer: Transformer-based model introduced in [xie2021segformer]
@@ IV-C1 Post-processing for Model Predictions
Common practices of evaluating deep fault delineation networks involve two steps [mustafa2024visual, LI2023105412, an2020overlap, an2021deep]: $(i)$ thresholding the network predictions to obtain binary outputs, and $(ii)$ comparing the binary outcomes with the fault test labels using a metric. The test labels are processed with the same morphological operations as the training faults in order to achieve consistent fault thickness at the input and output of a model. Consequently, the thresholding in step $(i)$ needs to be adaptive to the model and the data accordingly, followed by the same morphological operations for numerical evaluation. We compute the optimal threshold for the dataset using the Optimal Dataset Scale (ODS) metric [an2023understanding]. For a model, its optimal threshold is computed using the training set, which is then applied to binarize the predictions on the test data, followed by the same morphological operation for evaluation.
@@ IV-C2 Performance Metrics
As mentioned in Section II-E, pixel-based and distance-based metrics capture different aspects of prediction quality, and each can fail to fully reflect the structural accuracy of the predicted faults, making the evaluation of fault delineation methods an inherently challenging task. Our benchmark study thus adopts a holistic approach to assess prediction quality by considering a combination of multiple metrics alongside subjective inspection. In this study we choose to report one pixel-based metric, Dice Coefficient (defined in Eq. (1)) and two distance-based metrics: BCD and the modified Hausdorff (defined in Eq. 3, and Eq. 2, respectively). These metrics have not only been extensively used in the seismic literature for model performance assessment [wu2019faultSeg, guillon2020ground, 576361], but allow for two different evaluation axes: pixel overlap-based and structure-based.
@@ IV-C3 Fault Characteristic Comparisons
While the individual metrics described in Section II-F quantify dataset characteristics in isolation, comparative metrics evaluate the similarity between predicted and ground-truth faults. We describe below the metrics we use in Table VII to extend the individual descriptors into differences, ratios, and statistical comparisons.
@@ V-A2 Domain Shift and Joint Training
To further explore mitigation strategies for catastrophic forgetting, we conducted a focused case study using Elastic Weight Consolidation (EWC) with the deeplabv3 (ResNet-50 backbone) architecture, when transferring from the synthetic FaultSeg3D dataset to the real CRACKS and Thebe volumes. The EWC regularizer constrains the update of parameters that are important for the source task, aiming to preserve previously acquired knowledge during fine-tuning. As shown in Table IV, the expected behavior holds for the FaultSeg3D to CRACKS setting: performance on the target domain decreases slightly relative to standard fine-tuning, but performance on the source domain improves, indicating partial retention of source information. However, this pattern does not hold for the FaultSeg3D to Thebe case, where performance drops on both domains. This outcome is consistent with our broader observations that Thebe represents a strong outlier whose distribution is sufficiently distinct that attempts to retain source-domain information hinder learning of its features altogether. These results highlight that a driving factor in catastrophic forgetting is the distributional structure of the source-target pair.
The empirical pattern we observe (severe performance loss on the source domain after fine-tuning on Thebe, and the mixed effect of EWC across transfer pairs) can be explained by the interaction of three seismic-specific factors. First, faults are sparse, line-like signals: a network’s predictive capacity depends on a small number of localized, edge-sensitive features. Second, amplitude and contrast statistics differ markedly between datasets (e.g., Thebe’s standard deviation = 0.124 vs. CRACKS = 1.149 and FaultSeg3D = 1.052; see Fig. 3); such compression of dynamic range alters internal activation distributions (batchnorm and convolutional responses), rendering source-trained low-level filters less discriminative. Third, regularization schemes like EWC constrain parameters important for the source via a Fisher-based penalty: this helps when the target is moderately different (it preserves source knowledge while allowing limited adaptation, as in FaultSeg3D to CRACKS), but it can prevent learning altogether when the target is an extreme outlier (as in FaultSeg3D to Thebe), producing simultaneous degradation on both domains. Together these effects make seismic catastrophic forgetting both more likely and more abrupt than in many dense, object-centric vision tasks. Practical mitigations that follow from this analysis potentially include input-level alignment (histogram matching or style transfer), adaptive normalization layers, selective (layer-wise) fine-tuning, rehearsal or pseudo-replay, and capacity-aware choices (larger models or explicit adaptation for small models).
As an additional baseline, experiments on joint training on all possible combinations of datasets in a single training round are conducted using the unet model with a ResNet50 backbone. The results for all unet experiments, including these joint training configurations, are shown in Table II. The results further showcase that features learned from CRACKS and FaultSeg3D data align together, but these learned features are not easily transferable to Thebe. On the other hand, Thebe may act as a regularizer when paired with a large dataset like FaultSeg3D, hurting performance on seen datasets (compared to training them individually) but providing modest generalization to unseen ones: performance on the unseen dataset for this setting (i.e. CRACKS) does not drop as badly as that of the unseen dataset for other joint settings (e.g. FaultSeg3D + CRACKS).
As an additional step to assess the robustness and reliability of our results, we also perform our unet experiments on FaultSeg3D using 5 fold cross validation, which we report on Table III. We can observe that the results in these experiments deviate minimally from the ones we originally reported in Table II, showcasing that our experimental framework is robust against stochasticity.
TABLE II: Performance Metrics (DICE, BCD, Hausdorff) for different training schemes on Unet. All values are rounded to three decimals. Schemes are ranked as best, second best or worst based on satisfying at least two of the three metric criteria, to account for cases where metrics disagree. (blue for highest, pink for second, and gray for the worst.)
@@
Thebe 0.111 (+0.0347 , -0.0380) 18.767 (+33.9692 , -6.4945) 16.095 (+33.6256 , -6.3892) 0.020 (+0.0082 , -0.0077) 69.043 (+17.8047 , -13.3343) 49.012 (+19.7797 , -12.1662) 0.182 (+0.0230 , -0.0158) 52.054 (+7.3546 , -8.1523) 31.615 (+8.2126 , -4.7123)
FaultSeg3D 0.4165 (+0.0473, -0.0717) 9.5632 (+14.6540, -5.2962) 7.1908 (+12.4008, -4.1851) 0.1604 (+0.0479, -0.0445) 27.0161 (+8.7621, -6.1338) 20.6379 (+8.1430, -5.3172) 0.0417 (+0.0085, -0.0081) 232.7213 (+29.4637, -28.2944) 189.9385 (+27.9787, -26.0703)
Cracks 0.012(+0.057 , -0.008) 76.904 (+60.57 , -30.058) 57.798 (+60.57 , -30.058) 0.333 (+0.0574 , -0.0819) 11.330 (+10.8538 , -3.1426) 7.341 (+7.1827 , -2.4529) 0.001 (+0.0013 , -0.0007) 1377.945 (+550.5007 , -578.69) 1022.510 (+222.6350 , -321.1643)
Combined Training
FaultSeg3D + Cracks 0.662 (+0.115 , -0.082) 6.143 (+18.342 , -2.673) 5.282 (+15.365 , -1.890) 0.333 (+0.026 , -0.017) 15.436 (+4.649 , -1.583) 11.597 (+5.264 , -2.119) 0.001 (+0.0012 , -0.0007) 659.708 (+37.582 , -49.906) 430.887 (+31.481, -39.851)
FaultSeg3D + Thebe 0.060 (+0.025 , -0.018) 29.834 (+18.619 , -7.824) 20.151 (+14.253 , -5.342) 0.167 (+0.018 , -0.011) 29.717 (+7.782 , -3.947) 26.136 (+6.492 , -2.371) 0.211 (+0.012 , -0.008) 52.952 (+8.459 , -4.671) 43.348 (+7.935 , -3.649)
Cracks + Thebe 0.038 (+0.0475 , -0.0236) 34.183 (+17.2146 , -9.1272) 22.751 (+15.6639 , -7.0688) 0.195 (+0.039 , -0.0389) 17.8 (+6.04 , -3.61) 13.13 (+6.54 , -3.4) 0.112 (+0.021 , -0.0178) 144.67 (+35.17 , -24.26) 98.25 (+21.13 , -18.07)
All 0.337 (+0.2988 , -0.2512) 28.569 (+55.1384 , -20.0588) 23.674 (+35.9425 , -16.5400) 0.1899 (+0.032 , -0.04) 25.62 (+10.37 , -7.3) 17.88 (+10.7 , -7.3) 0.144 (+0.028 , -0.033) 89.69 (+55.91 , -25.90) 75.85 (+58.49 , -27.09)
Fine-Tuning
Cracks $\rightarrow$ FaultSeg3D 0.667 (+0.152 , -0.015) 4.994 (+20.04 , -2.603) 3.821 (+15.34 , -2.923) 0.028 (+0.0270, -0.0328) 73.433 (+19.3709, -15.0939) 45.378 (+16.5661, -12.8940) 0.027 (+0.0026, -0.0026) 165.315 (+14.13, -10.47) 99.996 (+10.03, -5.91)
Cracks $\rightarrow$ Thebe 0.099 (+0.02, -0.034) 21.959 (+34.82, -7.034) 15.771 (+10.93, -4.034) 0.159 (+0.0270, -0.0328) 26.070 (+8.2453, 6.2880) 20.929 (+7.4229, -5.2930) 0.181 (+0.0172, -0.0132) 53.808 (+11.815, -9.61) 42.169 (+13.56, -12.09)
FaultSeg3D $\rightarrow$ Cracks 0.0825 (+0.1060, -0.0587) 136.2934 (+256.0289, -72.6493) 134.0948 (+255.0063, -71.8545) 0.3751 (+0.0539, -0.0650) 8.8064 (+5.0173, -2.6473) 5.5773 (+3.4140, -2.0301) 0.0193 (+0.0055, -0.0045) 213.6949 (+21.8752, -18.4942) 147.2127 (+14.6880, -11.1500)
FaultSeg3D $\rightarrow$ Thebe 0.0745 (+0.1185, -0.0550) 93.4884 (+91.8717, -43.4160) 84.4202 (+84.7160, -38.6755) 0.1349 (+0.0246, -0.0233) 28.7014 (+7.7890, -6.2796) 23.3679 (+8.1318, -6.1587) 0.1582 (+0.0149, -0.0148) 78.6831 (+48.4932, -25.2388) 51.6490 (+47.2406, -19.6911)
Thebe $\rightarrow$ Cracks 0.019 (+0.0082 , -0.0077) 56.145 (+55.3901 , -18.078) 42.899 (+59.5264 , -15.8497) 0.317 (+0.0454 , -0.0514) 10.896 (+3.7560 , -2.1167) 6.656 (+3.2688 , -1.8525) 0.019 (+0.0052 , -0.0050) 161.131 (+14.8345 , -14.0856) 88.344 (+9.4150 , -9.1093)
Thebe $\rightarrow$ FaultSeg3D 0.526 (+0.1585 , -0.2580) 9.319 (+42.6775 , -6.5765) 8.182 (+40.3980 , -5.9642) 0.053 (+0.0116 , -0.0128) 38.141 (+9.1915 , -5.8445) 34.319 (+9.1316 , -5.7956) 0.037 (+0.0029 , -0.0027) 134.522 (+8.3162 , -6.5321) 122.936 (+8.0421 , -6.2061)
TABLE III: U-Net performance on FaultSeg3D dataset using 5-fold cross-validation
Metric Result
Dice $0.4493\pm 0.0285$
Hausdorff $6.1464\pm 0.5806$
BCD $8.6323\pm 0.7638$
TABLE IV: Results on EWC case study
Training configuration Source Target
Dice Hausdorff BCD Dice Hausdorff BCD
FaultSeg3D $\rightarrow$ Cracks 0.025 35.562 48.598 0.1998 7.5489 12.627
FaultSeg3D $\rightarrow$ Thebe 0.012 148.567 179.181 0.061 77.232 112.493
V-A3 Model Capacity and Transferability
It is generally established in the literature that pretraining on a large dataset can boost the performance of a model even in self-supervised settings [rs16050922]. However, in seismic applications, the benefits of pretraining are critically sensitive to the distributional similarity between the source and target domains—a phenomenon that reflects the strong coupling between geologic variability and model transferability. As shown in our experiments, models pretrained on FaultSeg3D or CRACKS and fine-tuned on Thebe underperform compared to models trained from scratch or initialized from ImageNet weights (Fig. 9). In such distributionally mismatched cases, the standard pretaining-finetuning strategy may hinder rather than help performance.
A factor that also plays a role in modulating generalization is model capacity. Fig. 9 shows that large models like segformer benefit from pretraining on CRACKS, while smaller models (e.g., hed, deeplab-m) generalize better when trained from scratch. The observation suggests that both data alignment and model capacity affect transfer effectiveness. Larger architectures, particularly transformer-based models like segformer, possess greater representational capacity and self-attention mechanisms that capture long-range spatial dependencies, an important property for preserving fault continuity and contextual consistency across sections. These characteristics can potentially explain their stronger cross-domain generalization when pretrained features are well aligned. However, the same capacity also makes them more sensitive to distributional mismatches, leading to potential overfitting to domain-specific amplitude statistics or noise patterns when source and target differ substantially. Smaller-capacity models, by contrast, may act as an implicit regularizer, limiting over-specialization and enabling better zero-shot transfer in mismatched scenarios, as can be observed in Fig. 9.
Furthermore, models respond differently to fine-tuning. When pretraining on FaultSeg3D and fine-tuning on Thebe, models degrade in CRACKS performance (Fig. 10, blue circle) shows that Thebe induces domain shifts that are hard to unlearn. Conversely, even though the finetuning dataset is distributionally closer to the target, Thebe-pretrained segformer and deeplab also degrade on CRACKS after finetuning on FaultSeg3D as shown in Fig.  10 (green circles). The asymmetric behavior highlights the difficulty of finding universally robust pretraining strategies.